OpenStudy (anonymous):

find dx/dy of x^2 + y^2 = 16

hartnn (hartnn):

do you know chain rule ?

hartnn (hartnn):

chain rule : d/dy (x^2) = 2x dx/dy

OpenStudy (anonymous):

the answer on the back of the book says -x/y. how do i get there?

hartnn (hartnn):

did you get how d/dy (x^2) = 2x dx/dy ??

OpenStudy (anonymous):

no

hartnn (hartnn):

ok, thats why i asked whether you already know the chain rule.

OpenStudy (anonymous):

2x+2y*y'=0

hartnn (hartnn):

if we take derivative of x^2 w.r.t x, then we get 2x but by chain rule, when we take derivative of x^2 w.r.t y we get x^2 dx/dy so we now have x^2 dx/dy + 2y = 0 isolate dx/dy from here.

OpenStudy (nincompoop):

@hartnn so this is essentially 2x * x' + 2y = 0 x' = -y/x

OpenStudy (anonymous):

I'm assuming that OP typed "dx/dy" as a typo. I think OP meant "dy/dx" as it corresponds to the answer in the back of the book, "-x/y".

OpenStudy (nincompoop):

dy/dx indicates with which respect to variable we are differentiating and in this case it's the y dx/dy would be the x

hartnn (hartnn):

dy/dx is -x/y yes x' = dx/dy = -y/x which gives dy/dx = -x/y too if the answer is -x/y then the question must be dy/dx